# Misspecification uncertainties in near-deterministic regression

### 02 Feb 2024 v1

@@ IV.2.2 Resampled POPS (rPOPS) ensemble
As shown in figure 3, the rPOPS results are excellent, giving tight agreement with test error distributions and only slightly more conservative errors bounds, a feature which is again robust across simulation engines (appendix C). A detailed investigation of the POPS-simplex and theoretical analysis of the rPOPS-ensemble is left as a topic for future study. We conclude with an application to challenging high-dimensional $P=\mathcal{O}(1000)$ datasets from atomic machine learning.
@@ IV.3 Interatomic potentials for material simulations
  • Empirical simulations with physically-motivated parametric models of interatomic interactions. These allow for large simulations over long timescales at a qualitative level of accuracy.
  • First-principle quantum simulations where interatomic interactions are obtained from approximate solutions of the full Schrodinger equation. Such calculations can provide quantitative predictions at a high computational cost that strongly limits the accessible time- and length-scales.
@@
The advent of machine learning has blurred the lines between these two categories by promising near-quantum accuracy at a much more favorable computational cost and scaling (typically cubic in the number of atoms for first-principles, and linear for empirical approaches) (Deringer et al., 2019). In the most common approaches, the total energy of the system of atoms $V$ is factored into a sum of per-atom contributions $V_{i}$, which are themselves parameterized in term of a vector of features that describe the local atomic environment around each atom ${\bf F}_{i}(\{{\bf x}_{j}|r_{ij}<r_{c}\})$, where $r_{c}$ is a interaction range. Training an ML model then proceeds by minimizing a squared error loss between the energies and forces (which are given by elements of the negative gradient of the total energy, $-\nabla V$) as obtained by first-principle methods and those predicted by the ML model. While certain ML approaches are in principle systematically improvable, the computational cost associated with evaluating energies and forces repeatedly for large systems and long times often limits the complexity of the models that are selected for applications. Further, quantum calculations can be tightly converged, so that their intrinsic error $\epsilon$ (w.r.t. a fully converged version of the same calculation) can be made small compared to typical errors incurred by the models that are selected in practice. This regime corresponds to the underparametrized deterministic limit considered above. Estimating the errors incurred by ML models of interatomic interactions (Wen and Tadmor, 2020; Li et al., 2018; Gabriel et al., 2021; Wan et al., 2021) is required when one wishes to assess the robustness of conclusions drawn from the simulations, especially when quantitative predictions are sought.
@@
In the following, training data was generated using an information-theoretic approach (Karabin and Perez, 2020; Montes de Oca Zapiain et al., 2022) that maximizes the information entropy of the training data distribution, as measured in feature space. 10,000 atomistic configurations containing between 2 and 32 tungsten atoms were generated and characterized using Density Functional Theory (Kohn and Sham, 1965), yielding 167,922 training data points in total (10,000 energies and 157,922 force components, 3 Cartesian for each atom in every configuration).
Per-atom energies are expressed as a linear combination of $P=1596$ so-called bispectrum components within the ‘Quadratic’ Spectral Neighborhood Analysis Potential (SNAP) formalism (Wood and Thompson, 2018), giving the linear surrogate model $V_{i}={\bm{\Theta}}\cdot{\bf F}_{i}(\{{\bf x}_{j}|r_{ij}<r_{c}\})$.
@@
Numerical experiments were carried out by randomly sub-selecting $N$ training points (either energies or gradient components), and by generating a discrete ensemble of $N$ pointwise-optimal models according to Eq. 21. The properties of the ensemble are then evaluated on a complementary set of 1,000 randomly-selected testing points (again either energies or gradient components). The random test/train partition is repeated 10 times for each value of $N/P$.
@@
This information thus generated is invaluable to estimate the uncertainties on physical quantities inferred from simulations that encounter local atomic environments that significantly differ from those contained in the training set. We however note that accurate uncertainty quantification relies on the training sets being representative of possible testing sets, which highlights the fundamental importance training data diversity (Karabin and Perez, 2020) and the perils of narrowly tailored training sets.
@@ V Conclusion
In spite of its ubiquity in practical applications, misspecification is often ignored in Bayesian regression approaches based on expected loss minimization, which can lead to erroneous conclusions regarding parameter uncertainty quantification, surrogate model selection, and error propagation. In this work, we show that the important near-deterministic underparameterized regime is amenable to approximation by a formally simple discrete ensemble ansatz that exploits the concentration of the cross-entropy minimizing ensemble onto pointwise optimal parameter sets. For linear regression problems, this ensemble can be obtained extremely efficiently using rank-one perturbations of the expected loss solution. POPS-constrained loss-minimizing ensembles are shown to produce excellent parameter distributions, worst-case bounds on model predictions, and approximations of the pointwise errors.
@@ Acknowledgements
Appendix A Model selection criteria for deterministic models
@@
We note that the value of the PAC-Bayes loss bound (8) satifies $BIC=2N\hat{\mathcal{L}}_{N}[\pi_{{\mathcal{L}},N}]-2C_{0}$, with $C_{0}$ model independent. In standard derivations ${\bm{\Sigma}}_{\mathcal{Y}}$ is set to the variational maximum likelihood solution ${\bm{\Sigma}}_{\mathcal{Y}}={\bm{\Sigma}}^{*}_{\mathcal{L}}$, yielding the familar expression $BIC=P\ln N+N\ln\|{\bm{\Sigma}}^{*}_{\mathcal{L}}\|+\ln\|{\bm{\mathcal{I}}}^{*}_{\mathcal{L}}\|$, where model independent terms are neglected. However, for deterministic regression problems, ${\bm{\Sigma}}_{\mathcal{Y}}$ is fixed. Neglecting model-independent terms, this gives an alternative BIC for deterministic models of
@@
Appendix B POPS-constrained variational ansatz for PAC-Bayesian linear regression
@@
However, the full variational scheme has comparable complexity to Bayesian ridge regression for linear models [Tipping, 2001]. Our variational ansatz replaces ${\bm{\Theta}}^{*}_{\mathcal{L}}$ in the pointwise fits (21) with a variational parameter ${\bm{\Theta}}_{V}$, giving a variational posterior mean and covariance of
@@
As for Bayesian ridge regression, the only non-linearity in (26) is the term ${\bm{\Sigma}}^{-1}_{V}({\bm{\Theta}}):{\bm{\Sigma}}^{\prime}_{V}({\bm{\Theta}}^{*}_{V})$, meaning any implementation can use existing iterative schemes e.g. [Tipping, 2001],[Wipf and Nagarajan, 2007]. A full study of the low data regime is left for future work.
Appendix C Additional plots of test error prediction

### 09 Apr 2024 v2

@@ IV.2.2 Resampled POPS (rPOPS) ensemble
As shown in figure 3, the rPOPS results are excellent, giving tight agreement with test error distributions and only slightly more conservative errors bounds, a feature which is again robust across simulation engines (appendix G). A detailed investigation of the POPS-simplex and theoretical analysis of the rPOPS-ensemble is left as a topic for future study. We conclude with an application to challenging high-dimensional $P=\mathcal{O}(1000)$ datasets from atomic machine learning.
@@ IV.3 Interatomic potentials for material simulations
  • Empirical simulations with physically-motivated parametric models of interatomic interactions. These allow for fast simulations at a qualitative level of accuracy.
  • First-principle quantum simulations obtained from approximate solutions of the full Schrodinger equation. These incur a high computational cost and exhibit poor scalability with system size.
@@
The advent of ML has blurred the lines between these two limits by promising near-quantum accuracy at a much more favorable computational cost and scaling (Deringer et al., 2019). The total energy of the system $V$ is typically factored into a sum of per-atom contributions $V_{i}$, which are themselves parameterized in term of a feature vector that describes the local atomic environment ${\bf F}_{i}(\{{\bf x}_{j}|r_{ij}<r_{c}\})$, where $r_{c}$ is a interaction range. Training an ML model then proceeds by minimizing a squared error loss between reference energies and gradients. While certain ML approaches are systematically improvable, the computational cost associated with repeated evaluations can limit the complexity of the models that are selected for applications. Further, quantum calculations can be tightly converged, so that their intrinsic error $\epsilon$ (w.r.t. a fully converged calculation) can be made small compared to typical errors incurred by computationally-efficient models. This regime corresponds to the underparametrized deterministic setting considered here. Estimating the errors incurred by such ML models is highly sought after to assess the robustness of conclusions drawn from the simulations (Wen and Tadmor, 2020; Li et al., 2018; Gabriel et al., 2021; Wan et al., 2021) .
@@
In the following, training data was generated using an information-theoretic approach (Karabin and Perez, 2020; Montes de Oca Zapiain et al., 2022). 10,000 atomistic configurations containing between 2 and 32 tungsten atoms were generated and characterized using Density Functional Theory (Kohn and Sham, 1965), yielding 167,922 training data points in total (10,000 energies and 157,922 gradient components). Per-atom energies are expressed as a linear combination of $P=1596$ so-called bispectrum components within the ‘Quadratic’ Spectral Neighborhood Analysis Potential (SNAP) formalism (Wood and Thompson, 2018), which can be expressed as a linear surrogate model $V_{i}={\bm{\Theta}}\cdot{\bf F}_{i}(\{{\bf x}_{j}|r_{ij}<r_{c}\})$.
@@
Numerical experiments were carried out by randomly sub-selecting $N$ training points (either energies or gradient components), and by generating a discrete ensemble of $N$ pointwise-optimal models according to Eq. 23. The properties of the ensemble are then evaluated on a complementary set of 1,000 testing points. The random test/train partition is repeated 10 times for each value of $N/P$.
@@ V Conclusion
stIn spite of its ubiquity in practical applications, misspecification is often ignored in Bayesian regression approaches based on expected loss minimization, which can lead to erroneous conclusions regarding parameter uncertainty quantification, surrogate model selection, and error propagation. In this work, we show that the important near-deterministic underparameterized regime is amenable to approximation by a formally simple discrete ensemble ansatz that exploits the concentration of the cross-entropy minimizing ensemble onto pointwise optimal parameter sets. For linear regression problems, this ensemble can be obtained extremely efficiently using rank-one perturbations of the expected loss solution. POPS-constrained loss-minimizing ensembles are shown to produce excellent parameter distributions, worst-case bounds on model predictions, and approximations of the pointwise errors.
@@ Acknowledgements
Appendix A PAC-Bayes concentration inequalities for regression
PAC-Bayes inequalities were originally designed for classification problems where the loss function is bounded to $[0,1]$. The extension to regression requires treatment of unbounded losses, which is discussed in detail in (Germain et al., 2016).
Determining the value of the upper bound for $C_{0}[\pi_{0}]$ in the PAC-Bayes concentration inequality (7) requires certain assumptions on the distribution of the error between the empirical and true expected loss (Germain et al., 2016). Following (Germain et al., 2016), we assume that this error distribution is unbounded but sub-normal, i.e. all error fluctuations are bounded by a normal distribution of variance $s^{2}$, giving $C_{0}[\pi_{0}]<s^{2}/2$. We refer the reader to (Germain et al., 2016) and subsequent works (Shalaeva et al., 2020; Masegosa, 2020) for further discussion.
Appendix B Model selection criteria for deterministic models
@@
We note that the value of the PAC-Bayes loss bound (7) satifies $BIC=2N\hat{\mathcal{L}}_{N}[\pi_{{\mathcal{L}},N}]-2C_{0}$, with $C_{0}$ model independent, as discussed in (Germain et al., 2016). In standard derivations ${\bm{\Sigma}}_{\mathcal{Y}}$ is set to the variational maximum likelihood solution ${\bm{\Sigma}}_{\mathcal{Y}}={\bm{\Sigma}}^{*}_{\mathcal{L}}$, yielding the familar expression $BIC=P\ln N+N\ln\|{\bm{\Sigma}}^{*}_{\mathcal{L}}\|+\ln\|{\bm{\mathcal{I}}}^{*}_{\mathcal{L}}\|$, where model independent terms are neglected. However, for deterministic regression problems, ${\bm{\Sigma}}_{\mathcal{Y}}$ is fixed. Neglecting model-independent terms, this gives an alternative BIC for deterministic models of
@@
Appendix C Summary of Laplace’s method
Laplace’s method, also known as the steepest descents method, is a well-known identity allowing the evaluation of an integral of an exponentiated function multiplied by a large number. The method applies to a function $f(x)$ which is twice differentiable and has a unique maximum in some closed interval $[a,b]$ which may be the entire real line. This encompasses all applications in this paper; the extension to multidimensional functions is straightforward. Laplace’s method then reads
$\lim_{N\to\infty}\int_{a}^{b}\exp[Nf(x)]{\rm d}{x}=\sqrt{\frac{2\pi}{N}}\sqrt{\frac{{\rm d}^{2}f((x)}{{\rm d}x^{2}}}^{-1}\exp[N\max_{x\in[a,b]}f(x)].$ (26)
Appendix D POPS-constrained variational minimization of ensemble ansatz
For finite $N/P$, we have access to training data ${\mathcal{D}}_{N}=\{{\mathbf{Y}}_{i},{\mathbf{X}}_{i}\}_{i=1}^{N}$ to give averages over $\langle\dots\rangle_{N}$, with some set of pointwise fits $\mathcal{T}_{E}=\{{\bm{\Theta}}_{i}\}_{i=1}^{N}$, ${\bm{\Theta}}_{i}\in{\mathcal{P}}_{\mathcal{M}}({\mathbf{X}}_{i})$. Our ensemble ansatz then reads
$\pi_{E}({\bm{\Theta}})=\langle{\rm w}_{E,i}\mathcal{N}({\bm{\Theta}}|{\bm{\Theta}}_{i},\sigma^{2}\mathbb{I})\rangle_{N}=\sum_{{\mathbf{X}}_{i}\in{\mathcal{D}}_{N}}{\rm w}_{E,i}{\rm w}_{i}\mathcal{N}({\bm{\Theta}}|{\bm{\Theta}}_{i},\sigma^{2}\mathbb{I}),$ (27)
where ${\rm w}_{E,i}$ are the discrete ensemble weights and ${\rm w}_{i}$ are the training data weights.
The empirical loss ^$\hat{\mathcal{L}}_{N}[\pi_{E}]$, which contributes All $\pi$-dependent terms in the PAC-Bayes upper bound (7) then read
^$\hat{\mathcal{L}}_{N}[\pi_{E}]=\int\hat{L}_{N}({\bm{\Theta}})\pi_{E}({\bm{\Theta}}),\quad KL(\pi_{E}|\pi_{0})=\int\ln\left|\frac{\pi_{E}({\bm{\Theta}})}{\pi_{0}({\bm{\Theta}})}\right|\pi_{E}({\bm{\Theta}}){\rm d}{\bm{\Theta}}.$ (28)
To proceed, we define a POPS-constrained variation $\delta_{i}{\bm{\Theta}}_{i}$ such that ${\bm{\Theta}}_{i}+\delta_{i}{\bm{\Theta}}_{i}\in\mathcal{P}_{\mathcal{M}}({\mathbf{X}}_{i})$, which gives a variation in $\pi_{E}$ of
$\delta_{i}\pi_{E}({\bm{\Theta}})\equiv\frac{{\rm w}_{E,i}{\rm w}_{i}}{\sigma^{2}}\delta{\bm{\Theta}}_{i}^{\top}\left[{\bm{\Theta}}-{\bm{\Theta}}_{i}\right]\mathcal{N}({\bm{\Theta}}|{\bm{\Theta}}_{i},\sigma^{2}\mathbb{I}).$ (29)
The upper bound (7) thus has a POPS-constrained variation
$\delta_{i}\left(\hat{\mathcal{L}}_{N}[\pi_{E}]+\frac{1}{N}KL(\pi_{E}|\pi_{0})\right)$ $=\frac{{\rm w}_{E,i}{\rm w}_{i}}{\sigma^{2}}\int\delta{\bm{\Theta}}_{i}^{\top}\left[{\bm{\Theta}}-{\bm{\Theta}}_{i}\right]\left(\hat{L}_{N}({\bm{\Theta}})+\frac{1}{N}\ln\left|\frac{\pi_{E}({\bm{\Theta}})}{\pi_{0}({\bm{\Theta}})}\right|\right)\mathcal{N}({\bm{\Theta}}|{\bm{\Theta}}_{i},\sigma^{2}\mathbb{I}){\rm d}{\bm{\Theta}},$
$={\rm w}_{E_{i}}{\rm w}_{i}\delta_{i}{\bm{\Theta}}_{i}^{\top}\int\nabla_{\bm{\Theta}}\left(\hat{L}_{N}({\bm{\Theta}})+\frac{1}{N}\ln\left|\frac{\pi_{E}({\bm{\Theta}})}{\pi_{0}({\bm{\Theta}})}\right|\right)\mathcal{N}({\bm{\Theta}}|{\bm{\Theta}}_{i},\sigma^{2}\mathbb{I}){\rm d}{\bm{\Theta}},$ (30)
where we use integration by parts. Aside from the POPS constrained variation, this is a standard variational result which in principle allows for variational minimization of $\pi_{E}({\bm{\Theta}})$ at finite $N$ for both linear and non-linear models. An alternative POPS-constrained variational approach for linear regression at finite $N$ is provided in appendix E.
In the limit $\sigma\to 0$ we make the approximation that, for ${\bm{\Theta}}-{\bm{\Theta}}_{i}=\delta{\bm{\Theta}}$, $\pi_{E}({\bm{\Theta}}_{i}+\delta{\bm{\Theta}})\simeq(1/N)\mathcal{N}(\delta{\bm{\Theta}}|{\bf 0},\sigma^{2}\mathbb{I})$. With a Gaussian prior of zero mean and covariance ${\bm{\Sigma}}_{0}$ we have
$\int\nabla_{\bm{\Theta}}\ln\left|\frac{\pi_{E}({\bm{\Theta}})}{\pi_{0}({\bm{\Theta}})}\right|\mathcal{N}({\bm{\Theta}}|{\bm{\Theta}}_{i},\sigma^{2}\mathbb{I}){\rm d}{\bm{\Theta}}=\int\left(\frac{{\bm{\Theta}}-{\bm{\Theta}}_{i}}{\sigma^{2}}+{\bm{\Sigma}}_{0}^{-1}{\bm{\Theta}}\right)\mathcal{N}({\bm{\Theta}}|{\bm{\Theta}}_{i},\sigma^{2}\mathbb{I}){\rm d}{\bm{\Theta}}\to{\bm{\Sigma}}_{0}^{-1}{\bm{\Theta}}_{i},\quad\sigma\to 0.$ (31)
In the present work we are primarily concerned with the underparametrized, or large data limit $P/N\to 0$, where the influence of the prior distribution is negligible, as demonstrated in section II.3. In this regime, for small $\sigma$ the variation of the upper bound (7) has a limiting form of
$\delta_{i}\hat{\mathcal{L}}_{N}[\pi_{E}]={\rm w}_{E_{i}}{\rm w}_{i}\delta_{i}{\bm{\Theta}}_{i}^{\top}\left[\nabla_{\bm{\Theta}}\hat{L}_{N}({\bm{\Theta}}_{i})+\frac{1}{N}{\bm{\Sigma}}_{0}^{-1}{\bm{\Theta}}_{i}\right]+\mathcal{O}(\sigma^{2}).$ (32)
This result shows the POPS-constrained variation of the PAC-Bayes loss upper bound converges to the gradient of the squared loss evaluated at a POPS-constrained parameter choice, to within a positive multiplicative constant . As a result, in the underparametrized limit the variational optimal choice of of POPS-constrained parameter ${\bm{\Theta}}_{\mathbf{X}}$ will be given by
${\bm{\Theta}}^{*}_{i}\equiv\arg\min_{{\bm{\Theta}}\in{\mathcal{P}}_{\mathcal{M}}({\mathbf{X}}_{i})}L({\bm{\Theta}}),$ (33)
in agreement with the result given in equation (21).
Appendix E POPS-constrained variational ansatz for PAC-Bayesian linear regression
@@
However, the full variational scheme has comparable complexity to Bayesian ridge regression for linear models (Tipping, 2001). Our variational ansatz replaces ${\bm{\Theta}}^{*}_{\mathcal{L}}$ in the pointwise fits (23) with a variational parameter ${\bm{\Theta}}_{V}$, giving a variational posterior mean and covariance of
@@
As for Bayesian ridge regression, the only non-linearity in (36) is the term ${\bm{\Sigma}}^{-1}_{V}({\bm{\Theta}}):{\bm{\Sigma}}^{\prime}_{V}({\bm{\Theta}}^{*}_{V})$, meaning any implementation can use existing iterative schemes e.g. (Tipping, 2001),(Wipf and Nagarajan, 2007). A full study of the low data regime is left for future work.
Appendix F Implementation of the ensemble ansatz in Python
The following python code, where A.T is the transpose of A and we assume a uniform weighting ${\rm w}_{E}({\bf X})=1$:
1import numpy as np
2def pointwise_optima(X,y):
3    # X.shape = (N,P), y.shape = (N,)
4    # Inverse of A
5    A_i = np.dot(X.T,X) +
6    # Minimum loss solution
7    T_L = np.linalg.solve(A_i,np.dot(X.T,y))
8    # Minimum loss errors
9    e = y - np.dot(X, T_L)
10    # Leverage of each point
11    X_dot_A = np.linalg.solve(A_i,X.T)
12    h = ( X * X_dot_A ).sum(1)
13    # Pointwise fits, shape = (P,N)
14    T = np.multiply(X_dot_A, e/h )
15    return T
An ensemble of test predictions can then be made simply by contracting with a constant tensor, a process which can easily be vectorized/parallelized.
Appendix G Additional plots of test error prediction